# MMSearch-Plus: Benchmarking Provenance-Aware Search for Multimodal Browsing Agents

### v2

@@ Appendix A Use of LLMs
The manuscript text was polished for fluency using Grammarly after the authors completed a full draft. The authors then manually reviewed all suggestions to ensure their appropriateness. In addition, ChatGPT was used to generate the task illustration in Figure 1 (upper left), based on the real image in Figure 8.
@@ Appendix C Comparison with Previous Benchmarks
| Benchmark/ Dataset | MM | Search |
| Text | Image |
| InfoSeek (Chen, 2023) | ✓ | ✓ | ✓ | ✗ | ✗ | ✗ | ✗ | ||||||
| FRAMES (Krishna et al., 2024) | ✗ | ✓ | ✗ | ✓ | ✓ | ✗ | ✗ | ||||||
| MMSearch (Jiang et al., 2024) | ✓ | ✓ | ✓ | ✗ | ✗ | ✗ | ✗ | ||||||
| FactualVQA(Wu et al., 2025) | ✓ | ✗ | ✓ | ✗ | ✗ | ✗ | ✗ | ||||||
| ChineseSimpleVQA (Gu et al., 2025) | ✓ | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ | ||||||
| BrowseComp (Wei et al., 2025) | ✗ | ✓ | ✗ | ✓ | ✓ | ✗ | ✗ | ||||||
| BrowseComp-ZH (Zhou et al., 2025) | ✗ | ✓ | ✗ | ✓ | ✓ | ✗ | ✗ | ||||||
| BrowseComp-VL (Geng et al., 2025) | ✓ | ✓ | ✓ | ✓ | ✓ | ✗ | ✗ | ||||||
| MM-BrowseComp (Li et al., 2025b) | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✗ | ||||||
| Ours | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ||||||
@@ Appendix D Related Work
A growing body of work argues that multimodality should go beyond tool orchestration to think with images. Methods such as V* (Wu & Xie, 2024) and instruction-guided masking (Zheng et al., 2024) improve visual grounding inside LLMs. RL-style training further enforces image-first reasoning: GRIT (Fan et al., 2025), DeepEyes (Zheng et al., 2025b), Pixel Reasoner (Su et al., 2025a), and Visionary-R1 (Xia et al., 2025) encourage explicit visual deliberation before action. Visual Agentic Reinforcement Fine-Tuning (Liu et al., 2025), OpenThinkIMG (Su et al., 2025b), and One-RL-to-See-Them-All (Ma et al., 2025) propose unified training frameworks for multimodal agents, while broader systems like PyVision (Zhao et al., 2025) and Thyme (Zhang et al., 2025) generalize the paradigm.
@@ Appendix G Evaluation Details
Appendix H Agent Behavior Analysis
@@
H.1 Episode Raster Plots
@@
H.2 Tool Transition Matrices
@@
H.3 Zoom-In Behavior Analysis
Figures 12 and 13 analyze zoom-in behavior patterns across agents. o3 shows the most aggressive zooming behavior, while Qwen and Gemini are more conservative.
Appendix I Rollout Trajectory Statistics
@@
Appendix J Prompts

### v3

@@ Appendix A Use of LLMs
The manuscript text was polished for fluency using Grammarly after the authors completed a full draft. The authors then manually reviewed all suggestions to ensure their appropriateness.
@@ Appendix C Comparison with Previous Benchmarks
| Benchmark/ Dataset | MM | Search | Multi-hop (2) | Thinking w/ Image | Multi- Image | ||
| Text | Image | Multi- turn | |||||
| InfoSeek (Chen, 2023) | ✓ | ✓ | ✓ | ✗ | ✗ | ✗ | ✗ |
| FRAMES (Krishna et al., 2024) | ✗ | ✓ | ✗ | ✓ | ✓ | ✗ | ✗ |
| MMSearch (Jiang et al., 2024) | ✓ | ✓ | ✓ | ✗ | ✗ | ✗ | ✗ |
| FactualVQA(Wu et al., 2025) | ✓ | ✗ | ✓ | ✗ | ✗ | ✗ | ✗ |
| ChineseSimpleVQA (Gu et al., 2025) | ✓ | ✗ | ✗ | ✗ | ✗ | ✗ | ✗ |
| BrowseComp (Wei et al., 2025) | ✗ | ✓ | ✗ | ✓ | ✓ | ✗ | ✗ |
| BrowseComp-ZH (Zhou et al., 2025) | ✗ | ✓ | ✗ | ✓ | ✓ | ✗ | ✗ |
| BrowseComp-VL (Geng et al., 2025) | ✓ | ✓ | ✓ | ✓ | ✓ | ✗ | ✗ |
| MM-BrowseComp (Li et al., 2025b) | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✗ |
| Ours | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ | ✓ |
@@ Appendix D Related Work
A growing body of work argues that multimodality should go beyond tool orchestration to think with images. Methods such as V* (Wu and Xie, 2024) and instruction-guided masking (Zheng et al., 2024) improve visual grounding inside LLMs. RL-style training further enforces image-first reasoning: GRIT (Fan et al., 2025), DeepEyes (Zheng et al., 2025b), Pixel Reasoner (Su et al., 2025a), and Visionary-R1 (Xia et al., 2025) encourage explicit visual deliberation before action. Visual Agentic Reinforcement Fine-Tuning (Liu et al., 2025), OpenThinkIMG (Su et al., 2025b), and One-RL-to-See-Them-All (Ma et al., 2025) propose unified training frameworks for multimodal agents, while broader systems like PyVision (Zhao et al., 2025) and Thyme (Zhang et al., 2025) generalize the paradigm.
@@ Appendix G Evaluation Details
Appendix H Results on MMSearch
To eliminate the impact of model’s internal knowledge, we obtain MMSearch from MMSearch+ by removing all tasks from MMSearch+ answerable without search by any model we tested. The results are reported in Table 4 and visualized in Figure 10 and 4.
In MMSearch, the best performing setting is full rollout with only text search (abbreviated as Text Search in the table) with o3 model, achieving 31.4% accuracy. This suggests that the current MLLMs cannot cleverly utilize fine-grained image search. In many categories, we observed a drop in accuracy from only text search to full rollout/full rollout + SoM. However, in certain categories like Sports, there is a great benefit of image search: for o3, the accuracy increases by 6.6 (full rollout) and 10.0 points (full rollout + SoM).
| Model / Search Mode | Avg | By Category | |||||||
| Geo. | Sports | Acad. | Film/TV | Tech | Games | Vlog | Music | ||
| Human | |||||||||
| Browser | 18.8 | 11.6 | 26.7 | 17.9 | 33.3 | 22.0 | 20.7 | 6.7 | 26.7 |
| Closed-source | |||||||||
| o3 (2025-04-16) | |||||||||
| Without Search | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
| Image Search | 10.0 | 14.0 | 13.3 | 10.3 | 16.7 | 4.9 | 17.2 | 3.3 | 0.0 |
| Text Search | 31.4 | 44.2 | 16.7 | 41.0 | 25.0 | 29.3 | 37.9 | 16.7 | 26.7 |
| Full Rollout | 27.3 | 46.5 | 33.3 | 17.9 | 8.3 | 14.6 | 37.9 | 16.7 | 33.3 |
| Full Rollout + SoM | 30.1 | 44.2 | 36.7 | 30.8 | 16.7 | 24.4 | 34.5 | 16.7 | 20.0 |
| GPT-5 | |||||||||
| Without Search | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
| Image Search | 8.8 | 9.3 | 13.3 | 7.7 | 25.0 | 4.9 | 13.8 | 3.3 | 0.0 |
| Full Rollout + SoM | 28.5 | 46.5 | 36.7 | 23.1 | 8.3 | 24.4 | 34.5 | 6.7 | 33.3 |
| Gemini-2.5-Pro | |||||||||
| Without Search | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
| Image Search | 10.5 | 14.0 | 13.3 | 12.8 | 25.0 | 7.3 | 10.3 | 3.3 | 0.0 |
| Full Rollout | 17.2 | 4.7 | 23.3 | 28.2 | 25.0 | 12.2 | 27.6 | 6.7 | 20.0 |
| Full Rollout + SoM | 20.9 | 23.3 | 13.3 | 28.2 | 25.0 | 19.5 | 24.1 | 20.0 | 6.7 |
| Open-source | |||||||||
| Qwen-2.5-VL-72B | |||||||||
| Without Search | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 | 0.0 |
| Image Search | 8.4 | 14.0 | 10.0 | 7.7 | 25.0 | 2.4 | 10.3 | 3.3 | 0.0 |
| Full Rollout | 3.3 | 4.7 | 3.3 | 0.0 | 8.3 | 4.9 | 3.4 | 3.3 | 0.0 |
| Full Rollout + SoM | 4.2 | 2.3 | 6.7 | 2.6 | 8.3 | 4.9 | 3.4 | 6.7 | 0.0 |
Appendix I Agent Behavior Analysis
@@
I.1 Episode Raster Plots
@@
I.2 Tool Transition Matrices
@@
In particular, Set-of-Mark allows image cropping and sub-image searching. Therefore, we investigate whether many of the cropped images were searched in the immediate next round, i.e., whether is large.
Gemini: 25.37% (34 out of 134 ZOOM actions)
Qwen: 10.56% (15 out of 142 ZOOM actions)
o3: 2.87% (18 out of 627 ZOOM actions)
I.3 Zoom-In Behavior Analysis
Figures 13 and 14 analyze zoom-in behavior patterns across agents. o3 shows the most aggressive zooming behavior, while Qwen and Gemini are more conservative.
Appendix J Rollout Trajectory Statistics
@@ I.3 Zoom-In Behavior Analysis
Appendix K Good Cases
Appendix L Prompts